Find the number of distinct integer 

solutions of the equation : x1+x2+ ......+x5=24

xi≥i and 1≤i≤5



1
Expert's answer
2021-08-02T17:24:12-0400

Let "y_i=x_i-i" for "1\leq i\leq5." We want to count solutions to


"y_1+y_2+y_3+y_4+y_5=24-1-2-3-4-5=9"

There are "C(9+ 5 - 1, 5 - 1)" ways to place "5" indistinguishable objects into "9" distinguishable boxes


"\dbinom{13}{4}=\dfrac{13!}{4!(13-4)!}=715"

There are 715 solutions.